***** All of the answers are correct. I just want to know How to calculate the ( Sales Volume (units) ) *****

Problem 12-4 (book/static) Question Help 0 Manitowoc Crane (B). Manitowoc Crane (U.S.) exports heavy crane equipment to severA
Answer #1

Hint given in Question:-

sales volume year 1 =10000 units

Sales Volume will increase at 12% per annum through year eight

Solution:-

Year Sales Volume units (calculation) Sales units
Year 1 10000 10000(given)
Year 2

10000+12%of 10000

10000 + 1200 = 11200

11200
Year 3

11200 + 12% of 11200

11200 + 1344 = 12544

12544
Year 4

12544 + 12% of 12544

12544 + 1505.28 = 14049.28

= 14049

14049
Year 5

14049 + 12% of 14049

14049 + 1685.88 = 15734.88

= 15735

15735
Year 6

15735 + 12% of 15735

15735 + 1888.2 = 17623.2

   = 17623

17623
Year 7

17623 + 12% of 17623

17623 + 2114.76 = 19737.76

= 19738

19738
Year 8

19738 + 12% of 19738

19738 + 2368.56 = 22106.56

   = 22107

22107